# Learning Hidden Physics and System Parameters with Deep Operator Networks

### v1

@@ Acknowledgments
VK and BP would like to acknowledge the computing support provided by The Deep Learning and GPU computing Platform (BD/DPA-BDP1) at Robert Bosch, GmbH, Renningen, Germany. DRS and SG would like to acknowledge computing support provided by the Advanced Research Computing at Hopkins (ARCH) core facility at Johns Hopkins University and the Rockfish cluster. ARCH core facility (rockfish.jhu.edu) is supported by the National Science Foundation (NSF) grant number OAC1920103. DRS and SG would like to acknowledge Dr. Krishna Kumar, Dr. Xingjian Li, and Dr. Hassan Iqbal for the insightful discussions during the journal club meetings.
@@
Writing—original draft: VK, DRS, BP, SG
Writing—review & editing: VK, DRS, BP, SG
@@ Appendix A Hyperparameter tuning
In the deep learning community, Weights & Biases (WandB) is a popular online platform that provides tools for monitoring and controlling hyperparameters during model training runs. Sweeps, a crucial component of WandB, allow for the systematic tuning of model hyperparameters. The Sweeps configuration dictionary contains the predefined hyperparameter search space. Depending on this setup, a sweep could include several runs, with grid search, random search, or Bayesian optimization used for hyperparameter sampling. Sarkar et al. [46] showcases the use of WandB for hyperparameter tuning on PINNs training. We have utilized the WandB tool for hyperparameter tuning of our framework for the parameter identification problem of the Burgers equation. The report of the hyperparameter tuning can be found at the link https://api.wandb.ai/links/droysar1-johns-hopkins-university/pzs4hfup.
@@
[…]
Branch -- ReLU
Trunk -- Tanh
MLP -- Tanh
10k $41e-$ 4 10
@@
Branch -- ReLU
Trunk -- Tanh
MLP -- Tanh
10k $41e-$ 4 1
@@
Figure 1: Loss function over the training process of optimal model: (a) Reaction diffusion equation with $N_{train}=500,N_{d}=500.$ (b) Burgers’ equation with $N_{train}=1000,N_{d}=500.$

### v2

@@ Acknowledgments
VK and BP would like to acknowledge the computing support provided by The Deep Learning and GPU computing Platform (BD/DPA-BDP1) at Robert Bosch, GmbH, Renningen, Germany. DRS and SG would like to acknowledge computing support provided by the Advanced Research Computing at Hopkins (ARCH) core facility at Johns Hopkins University and the Rockfish cluster. ARCH core facility (rockfish.jhu.edu) is supported by the National Science Foundation (NSF) grant number OAC1920103. DRS and SG would like to acknowledge Dr. Krishna Kumar, Dr. Xingjian Li, and Dr. Hassan Iqbal for the insightful discussions during the journal club meetings. DRS and SG were partially supported by JHU Discovery award 2024 and NSF grant number 2436738.
@@
Writing - original draft: VK, DRS
Writing - review & editing: VK, DRS, BP, SG
@@ Appendix A Hyperparameter tuning
In the deep learning community, Weights & Biases (WandB) is a popular online platform that provides tools for monitoring and controlling hyperparameters during model training runs. Sweeps, a crucial component of WandB, allow for the systematic tuning of model hyperparameters. The Sweeps configuration dictionary contains the predefined hyperparameter search space. Depending on this setup, a sweep could include several runs, with grid search, random search, or Bayesian optimization used for hyperparameter sampling. Sarkar et al. Sarkar et al. [2025a] showcases the use of WandB for hyperparameter tuning on PINNs training. We have utilized the WandB tool for hyperparameter tuning of our framework for the parameter identification problem of the Burgers’ equation. The report of the hyperparameter tuning can be found at the link https://api.wandb.ai/links/droysar1-johns-hopkins-university/pzs4hfup.
@@
Reaction-Diffusion $[101,128,128,128,50]$ $[2,128,128,128,50]$ $[3,128,128,128,1]$
Branch - ReLU
Trunk - Tanh
MLP - Tanh
10k 1e41e-4 10
@@
Branch - ReLU
Trunk - Tanh
MLP - Tanh
10k 1e41e-4 1
Heat Equation (2D)
CNN: $33\times 33$,
Channels $[1,40,60]$,
Kernel $3\times 3$, stride 22,
Maxpool $2\times 2$
FCN: $[240,64,64,64]$
$[3,128,128,128,64]$ $[5,128,128,128,1]$
Branch - ReLU, Linear
Trunk - Tanh, Linear
MLP - Tanh, Linear
2000 1e41e-4 1
@@
Heat $[440,64,64,64,100]$ $[3,64,64,64,100]$ $[440,32,16,8,1]$
Branch-Tanh
Trunk-Tanh
MLP - GeLU
80k
1e31e-3
exponential,
rate = 0.9
$75$
Helmholtz $[1024,32,32,32,100]$ $[2,32,32,32,100]$ $[1024,32,32,32,3]$
Branch-Tanh
Trunk-Tanh
MLP - GeLU
80k
1e31e-3
exponential,
rate = 0.9
$75$
Figure 1: Loss function over the training process of optimal model: (a) Reaction diffusion equation with $N_{train}=500,N_{d}=500.$ (b) Burgers’ equation with $N_{train}=1000,N_{d}=500.$ (c) Heat equation with $N_{train}=500,N_{d}=550.$
Figure 2: Sample 1: hidden physics solution comparison, mean absolute error = 0.03177
Figure 3: Sample 2: hidden physics solution comparison, mean absolute error = 0.03415
(a) Sample 1: solution field accuracy comparison, mean absolute error $u_{pred}(t\leq 1)=0.00507,u_{pred}(t>1)=0.04674,\\ u_{hpm}(t\leq 1)=0.00627,u_{hpm}(t>1)=0.03299$
(b) Sample 2: solution field accuracy comparison, mean absolute error $u_{pred}(t\leq 1)=0.06620,u_{pred}(t>1)=0.13954,\\ u_{hpm}(t\leq 1)=0.08864,u_{hpm}(t>1)=0.35198$
Figure 4: Reaction diffusion equation: Comparison of DHPO solution & HPM integrated solution for two representative test samples with input function from GRF function space $l=0.2$ (nonzero at boundaries). It should be noted that training data is collected from t = 0 to 1., extrapolation time region separated by the white horizontal lines.
(a) Sample 1: solution field accuracy comparison, mean absolute error $u_{pred}(t\leq 1)=0.00458,u_{pred}(t>1)=0.05970,\\ u_{hpm}(t\leq 1)=0.0059,u_{hpm}(t>1)=0.03579$
(b) Sample 2: solution field accuracy comparison, mean absolute error $u_{pred}(t\leq 1)=0.02584,u_{pred}(t>1)=0.12762,\\ u_{hpm}(t\leq 1)=0.02967,u_{hpm}(t>1)=0.13450$
Figure 5: Reaction diffusion equation: Comparison of DHPO solution & HPM integrated solution for two representative test samples with input function from Modified GRF function space $l=0.2$ (zero at boundaries). It should be noted that training data is collected from t = 0 to 1., extrapolation time region separated by the white horizontal lines.
(a) Sample 1: hidden physics solution comparison, mean absolute error = 0.00567.
(b) Sample 2: hidden physics solution comparison, mean absolute error = 0.01787.
Figure 6: Reaction diffusion equation: Comparison of reference and predicted PDE terms $\mathcal{N}$ with input function from sine function space. It should be noted that, the Reference $\mathcal{N}$ is computed using finite difference with the gradients being computed using second order scheme.
(a) Sample 1: hidden physics solution comparison, mean absolute error = 0.01587.
(b) Sample 2: hidden physics solution comparison, mean absolute error = 0.21618.
Figure 7: Reaction diffusion equation: Comparison of reference and predicted PDE terms with input function from GRF function.
(a) Sample 1: hidden physics solution comparison, mean absolute error = 0.01435.
(b) Sample 2: hidden physics solution comparison, mean absolute error = 0.07512.
Figure 8: Reaction diffusion equation: Comparison of reference and predicted PDE terms with input function from Modified GRF function.
(a) Sample 1: hidden physics solution comparison, mean absolute error = 0.01267.
(b) Sample 2: hidden physics solution comparison, mean absolute error = 0.05602.
Figure 9: Burgers’ equation: Comparison of reference and predicted PDE terms.
Figure 10: DeepONet solution: Sparse training solution comparison. mean absolute error at training time instants $(t=0,1)$ is 0.03029, mean absolute error at testing time instants $(0<t<1)$ is 0.38671.